OpenStudy (anonymous):

simplify 2x^2 + 3x +1 / 2x+1 and explain why f(1) = 2 and f(0) = 1 and f(-1) = 0 yet f(-1/2) is undefined. Please help.

OpenStudy (anonymous):

2x - {3x - 2 [ x - (1 - x)]} 2x - {3x - 2 [x - 1 + x] 2x - {3x - 2x + 2 - 2x} 2x - 3x + 2x - 2 + 2x 3x - 2

OpenStudy (anonymous):

(2x+1)(x+1)/(2x+1)

OpenStudy (anonymous):

f(1)=(2+1)(1+1)/(2+1)=2 f(0)=(1)(1)/(1)=1 f(-1)=(1-2)(1-1)/(1-2)=0 f(1/2)=(1-1)(1/2-1)/(1-1)=0/0 0/0 is undefined

OpenStudy (anonymous):

Wow thank you!! So (2x+1)(x+1)/(2x+1) is the simplest form?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[y=x+1, x \neq 1/2\] is simplest form

OpenStudy (anonymous):

Thank you so much :) and -1/2 is undefined because it equals 0/0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're the best!

OpenStudy (anonymous):

all in a days work :^)

OpenStudy (anonymous):

:D